OpenStudy (anonymous):

What is the slope of the line whose equation is 2x – 4y = 10?

OpenStudy (anonymous):

First thing you should do is put it in y= form. do you know how to do that?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

what y= form is, is a form where you can plug in a coordinate for x to find y

OpenStudy (anonymous):

y=ax+b

OpenStudy (anonymous):

The first step is to get y alone. DO you know how to do that?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

2x – 4y = 10 first subtract 2x from each side 2x – 4y = 10 -2x -2x to get – 4y = 10-2x

OpenStudy (anonymous):

Now you want get rid of that -4 in front of the y

OpenStudy (anonymous):

To do that you divide everything by -4 – 4y = 10-2x /-4 /-4 /-4 to get y=-2.5+.5x

OpenStudy (anonymous):

sssssssssoooooooooooooooo what's the answer mate

OpenStudy (anonymous):

to make it easier you can switch -2.5 and .5x around y=.5x-2.5

OpenStudy (anonymous):

that's y intercept from

OpenStudy (anonymous):

I assume you know how to find the slope from here

OpenStudy (anonymous):

no, no I do not

OpenStudy (anonymous):

Do you know what the slope of a line is?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

The slope is the steepness of a line

OpenStudy (anonymous):

in this equation if -2.5 represents where the line crosses the y intercept and y and x are coordinated that plug in, what can we conclude the slope is?

OpenStudy (anonymous):

I have no idea, can you just tell me ma'am I do not understand

OpenStudy (anonymous):

@aszarek8999

OpenStudy (anonymous):

https://www.desmos.com/calculator

OpenStudy (anonymous):

plug in your equation here

OpenStudy (anonymous):

remember the slope is the steepness of the line

OpenStudy (anonymous):

in y intercept form the slope in plainly stated in the equation. No math required